# Prevention is better than cure? Feedback from high specific energy winds in cosmological simulations with Arkenstone

### v1

@@ 3.2.3 Stellar Mass Black Hole Mass Relation
Figure 6: BH-stellar mass relation at $z=0$. Line and points are displayed in the same way as Fig. 5, with a fit to observational data from Kormendy & Ho (2013) shown as a dotted line. BHs in Arkenstone tend to be slightly more massive than in TNG by $z=0$, likely causing changes in quenching behaviour due to TNG’s feedback model.
@@
Though we have repeatedly discussed the caveats associated with the complex interplay between Arkenstone and the fiducial TNG BH model, particularly for galaxies at higher masses, it is still worth looking at how the BH population evolves in these simulations. Therefore, in Fig. 6 we show the stellar mass - black hole mass relation for our simulations. This was another of the observationally motivated relations used to calibrate the TNG model. The inclusion of MHD increases this quantity slightly in the TNG framework, so that the full fiducial TNG model lies slightly above the Kormendy & Ho (2013) relation (see Pillepich et al., 2018a).
@@
In most Arkenstone runs, BHs are a factor of a few larger than the equivalent in TNG. Investigating the reasons why this happens is beyond the scope of this paper, and will be studied in future work. But we note that none of the parameters for the AGN feedback model used in TNG were changed, despite a dramatic change in how galactic winds work. The larger BHs do have implications for galaxy quenching, however, due to the activation of TNG’s powerful kinetic feedback mode having a mass dependent switch (and typically activates in galaxies with stellar masses above of $10^{10.5}\,$, Weinberger et al., 2017).
@@
The only discernible difference in the stellar mass - BH mass relation in our run variations lies at the massive end of the $\eta_{\mathrm{E}}=0.9$ run, whose black holes tend to lie below the Kormendy & Ho (2013) relation and the other Arkenstone runs. This implies that highly energy loaded winds can suppress the growth of black holes in large haloes, though we defer a full investigation of Arkenstone in group and cluster-scale haloes to future work.
@@ 3.2.4 Galaxy Sizes
Figure 7: Galaxy sizes at $z=0$, defined as the stellar half mass radii, as a function of stellar mass. Lines and points are displayed in the same way as Fig. 5, with observational data from van der Wel et al. (2024) included as dotted lines with shaded uncertainty areas. Arkenstone tends to produce smaller low-mass galaxies, with the growth of larger galaxies happening at slightly smaller masses than in TNG.
Interestingly, we find that the energy loading of our galactic winds plays a strong role in the stellar sizes of galaxies at $z=0$. In Fig. 7, we show the median 3D stellar half-mass radius as a function of galaxy stellar mass. TNG was calibrated to this metric, though to a different observational dataset than the one presented here. The TNG runs (solid black line) slightly over-estimate the size of galaxies compared to the latest observational data from JWST (van der Wel et al., 2024), though we note that MHD (as present in the original TNG simulations) tends to slightly reduce the size of stellar discs (Pillepich et al., 2018a). It is also worth noting that the TNG model significantly improved on the over-estimation of galaxy radii found in the original Illustris simulation (Genel et al., 2018).
With Arkenstone, the sizes of galaxies at a given stellar mass has a strong dependence on $\eta_{\mathrm{E}}$, with higher energy loadings leading to larger galaxies at nearly all galaxy masses (shown in the top panel of Fig. 7). The central value of $\eta_{\mathrm{E}}=0.3$ has the best match to the observational galaxy sizes at the low mass end, though we note that the turning point where the slope steepens happens at much lower masses in Arkenstone compared to both TNG and observations. This is likely linked to the galaxy becoming quenched by the TNG BH model’s kinetic feedback, which kicks in at slightly lower masses in the Arkenstone runs due to larger BH masses. We postpone a more detailed study of the structure of galaxies to future work. The bottom two panels show that galaxy sizes are essentially independent of mass loading, and also change little when including a variable energy loading (with perhaps a slight decrease in galaxy sizes at the massive end, though with few haloes this trend is uncertain).
@@ 3.3 CGM properties at $z=0$
Figure 9: Mass-weighted mean circumgalactic medium temperatures within $R_{200}$ at $z=0$, normalised by the nominal virial temperature $T_{200}$. Lines and points are displayed in the same way as in Fig. 5. CGM temperatures as a function of halo mass tend to have a “U”-shape, with a minimum temperature (relative to $T_{200}$) at halo masses of $\sim\!10^{11-11.5}\,$. Temperatures in Arkenstone runs tend to lie above those of TNG at $z=0$.
@@
Figure 10: Galaxy SFR as a function of stellar mass at $z=0$ for TNG (top) and the intermediate variable energy loading Arkenstone run ($\eta_{\rm E}\propto M_{\rm h}^{-1/3}$, bottom), coloured by the mass-weighted average gas fraction in the CGM (between twice the stellar half mass radius and $R_{200}$), normalised by the universal baryon fraction. The separated axes at the bottom of each panel show the galaxies with a zero SFR. Observational comparison data is from Alsing et al. (2024), with the median shown by a dotted black line and the inner and outer shaded regions corresponding to the $16-84$ and $2.5-97.5$ percentile ranges, respectively. With Arkenstone, the SFR in low mass galaxies is correlated with the gas fraction in the CGM – indicative of preventative feedback.
@@
In Fig. 8 we show the median gas fraction within $R_{200}$, relative to the cosmic baryon fraction, as a function of halo mass. The shape of the TNG line is set by the combination of the stellar and BH feedback models. TNG was also calibrated to this quantity, though data exists only for groups (of which there are only a handful in our box) and clusters (of which there are none). As this paper focuses on haloes below this scale, where no observational data exists, we do not show comparison data here.
@@
Below a halo mass of $\sim\!2\times 10^{12}$ M, the amount of gas in the CGM is regulated by stellar feedback. We have already discussed how TNG tends to rely on ejection of mass from the ISM to regulate star formation. We can see how this ejection also removes much of the CGM at the lowest halo masses, while slightly larger haloes can retain more material. Above $\sim\!2\times 10^{12}$ M, the kinetic mode of TNG’s AGN feedback typically kicks in, expelling much of the inner CGM. As the halo mass increases further, $R_{200}$ increases and haloes can gravitationally hold onto more of that material, leading to an increase in the gas fraction once more.
@@
A number of differences are present between runs with different energy loadings (top panel), with opposing trends either side of a halo mass of $\sim\!10^{12}$ M. At low masses, the higher energy loadings lead to more preventative feedback, lowering the gas fraction that ends up within $R_{200}$. At high masses, the highest energy loading actually leads to the retention of more gas. This is likely due to the stunted growth of black holes (see Fig. 6) leading to less feedback, particularly because fewer BHs will then enter TNG’s powerful, mass-dependent kinetic mode. Again, the mass loading of Arkenstone winds makes significantly less difference to the fraction of gas within the CGM compared to energy loading. The winds with variable energy loading smooth the abrupt change in median gas fraction present in TNG at halo masses of $\sim\!10^{10.5}$ M, and then retain a similar structure at high masses as the AGN feedback acts in the same way. Gas fractions within $R_{200}$ at $z=0$ in Arkenstone tend to lie below those of TNG – this CGM depletion can be seen in the regions close to galaxies in Fig. 3.
@@
Preventative feedback can act through the heating of the CGM (and IGM) around galaxies, inhibiting the cooling of gas onto the central galaxy. To study this, in Fig. 9 we show the mass-weighted temperature of all non-star forming gas cells (those not on the eEoS described in Section 2) within $R_{200}$, normalised by the halo virial temperature $T_{200}={\mu m_{\mathrm{p}}GM_{200}}/{(2k_{\mathrm{B}}R_{200})}$, where $\mu=0.59$ is the mean molecular weight of ionised primordial gas.
@@
The vast majority of haloes in all of the runs in Fig. 9 lie below 1, indicating a sub-virial mean temperature. TNG has CGM temperatures closest to the virial temperature at the lowest and highest masses, with a dip in between at $\sim\!10^{11-11.5}\,$. This dip likely corresponds to haloes with a virial temperature near the peak of the cooling curve, leading to more efficient cooling (Fielding et al., 2017), particularly in TNG where the CGM density is higher than in Arkenstone (see Fig. 8) Like the gas fraction plot, the relationship at low masses changes more smoothly in Arkenstone compared to TNG. The Arkenstone runs have a somewhat similar shape, with a minimum in CGM temperature in the halo mass range $\sim\!10^{11-11.5}\,$, but with the average $z=0$ temperature being higher. There are no clear changes between the different Arkenstone simulations, with the exception of hotter CGM temperatures for the lowest mass galaxies when fixed energy loadings are increased. We note that this median value of the mass-weighted mean temperatures of haloes hides significant complexity in the CGM, which generally cannot be characterised by a single virial temperature (Lochhaas et al., 2021).
@@
In Fig. 10 we see how this change in CGM properties drives the regulation of star formation. We show two-dimensional histograms of the SFR of galaxies as a function of stellar mass, highlighting the star-forming main sequence for TNG (top panel) and the central variable energy loading Arkenstone run (bottom panel). The histogram is coloured by the mass-weighted average CGM gas fraction (defined as the gas within $R_{200}$ minus the gas within twice the stellar half mass radius), normalised by the cosmic baryon fraction. The darker regions therefore indicate where the CGM is relatively depleted of gas, which for Arkenstone is interestingly correlated with lower SFR for low mass galaxies. This trend is not present in TNG, which tends to have larger CGM gas fractions in general. The combination of these highlights the preventative nature of feedback with high specific energy winds.
3.4 Preventative feedback in action at $z=2$
@@
In this section we show a subset of the above results at $z=2$, where the star formation rate is higher, and so the subsequent feedback is stronger. This allows us to illuminate the differences between how Arkenstone and TNG regulate star formation, even though they both can lead to similar $z=0$ galaxy properties. Note that in this section our plots also include the simulation with fixed $\eta_{\rm M}=0.9$ and $\eta_{\rm E}=0.3$, which did not reach $z=0$ and thus was not shown previously.
@@
In Fig. 11 we show the SMHM at $z=2$, which shows mostly similar trends to $z=0$. Changing the energy loading of the hot wind has the greatest effect on both the normalisation and (at low and intermediate halo masses) the slope of the SMHM (top panel), with a larger effect shown at smaller halo masses. The mass loading affects the normalisation of the SMHM but not the slope (central panel), with higher stellar masses with higher mass loadings. As discussed earlier, this relationship is due to lower mass loading causing a hotter and faster wind (with the same total power) and increased preventative feedback. The variable energy loadings (bottom panel) counter the overproduction of stars at the low mass end, bringing the SMHM closer to that of TNG. Notably, all of the simulations lie above the observationally inferred SMHM at $z=2$ from Shuntov et al. (2022).
@@
The most significant differences between the runs, which hold the key to understanding how Arkenstone differs from TNG, are shown in the gas fraction (Fig. 12) and temperature (Fig. 13) plots.
@@
Turning first to the gas fractions, for TNG this starts off low at the smallest halo masses, before increasing to $\sim\!80$ per cent of the cosmic baryon fraction at a halo mass of $10^{12}$ M, and slowly declining beyond that as the black hole feedback takes effect. For Arkenstone  the gas fraction is much flatter across halo masses. For simulations with different constant energy loading values, the normalisation for haloes below $10^{12}$ M is inversely proportional to $\eta_{\mathrm{E}}$. At the TNG ‘turnover’ mass of $10^{12}$ M, the Arkenstone runs have a gas fraction of 40-50 per cent of baryons, around half of TNG’s value at this point. The strong implication of this is that either high specific energy winds push gas outside $R_{200}$, or gas is prevented from accreting within $R_{200}$ in the first place. This ‘prevention-via-CGM-depletion’ scenario, in contrast to ISM ejection, is a central feature of the Arkenstone model. Notably, the relationship between energy loading and gas fraction reverses in haloes above $10^{12}$ M, though this is likely due to the variations in stellar feedback affecting black hole growth and feedback.
@@
As we have seen for many of the quantities in this paper, varying the mass loading affects the gas fractions, but has a much smaller impact than variations in energy loading. The lowest mass loading of $\eta_{\mathrm{E}}=0.1$ has also the lowest gas fractions for low and intermediate mass galaxies, indicating that hotter and faster winds reduce the baryon content of the CGM even more at higher redshift. The variable energy loadings on the other hand lead to a gradually increasing gas fraction as a function of halo mass, up to a halo mass of $\sim\!10^{12}$ M. This slope, and the peak gas fraction value of $\sim\!50$ per cent, are both much smaller than in TNG.
@@
The temperature of the CGM is also an indicator of preventative feedback, with hotter, diffuse gas having longer cooling times and so reducing inflows onto the central galaxy. The CGM temperature at $z=2$ is shown in Fig. 13, defined in the same way as Fig. 9. Here we again see a strong difference between TNG and Arkenstone. The CGM temperatures at the lowest masses considered are largely similar between the runs, but while TNG’s decreases to much cooler temperatures for haloes around $\sim\!10^{11}\,$, the Arkenstone runs stay hotter throughout the entire halo mass range.
@@
The temperature difference varies with energy loading (top panel), particularly in the halo mass range $10^{11-12}$ M. Higher energy loading leads to a hotter and more diffuse CGM, more able to prevent gas accreting onto galaxies and forming stars. The highest mass loaded Arkenstone run leads to a cooler CGM, as more energy is used to transport a heavier wind. The slope of the variable energy loading does not strongly impact the CGM temperature at $z=2$.
Across all Arkenstone runs, the median CGM temperature at $z=2$ is nearly always above that of TNG, even in the case of higher mass loadings. Across the parameter space studied in this work, preventative feedback therefore plays much more of a role, with the CGM being consistently hotter and more diffuse.

### v2

@@ 3.2.3 Stellar mass – black hole mass relation
Figure 6: BH-stellar mass relation at $z=0$. Line and points are displayed in the same way as Fig. 5. A fit to observational data from Kormendy & Ho (2013) is shown as a dotted line, with the grey shaded region showing the 0.29 dex intrinsic scatter about this relation. BHs in Arkenstone tend to be slightly more massive than in TNG by $z=0$, likely causing changes in quenching behaviour due to TNG’s feedback model.
@@
Though we have repeatedly discussed the caveats associated with the complex interplay between Arkenstone and the fiducial TNG BH model, particularly for galaxies at higher masses, it is still worth examining basic galaxy versus BH scaling relations for the Arkenstone models. In Fig. 6 we show the stellar mass - black hole mass relation for our simulations. This was another of the observationally based relations used to calibrate the TNG model. The inclusion of MHD increases this quantity at fixed stellar mass slightly in the TNG framework, so that the full fiducial TNG model lies slightly above the median Kormendy & Ho (2013) relation (see Pillepich et al., 2018a).
@@
In most Arkenstone runs, BHs are a factor of a few larger than the equivalent in TNG. Investigating the reasons why this happens is beyond the scope of this paper, and will be studied in future work. But we note that none of the parameters for the AGN feedback model used in TNG were changed for the Arkenstone runs, despite a dramatic change the stellar driven wind properties. The larger BHs do have implications for galaxy quenching, however, due to the BH mass dependent switch that determines when TNG’s powerful kinetic feedback mode is activated (typically in galaxies with stellar masses above of $10^{10.5}\,$M, Weinberger et al., 2017).
@@
There appears to be a slight dependence of the amplitude of the stellar mass – black hole mass relation on the energy loading of the wind, with higher energy loadings leading to slightly more massive black holes in the dwarf regime, and slightly less massive black holes in the group regime. This implies that highly energy loaded winds can potentially suppress the growth of massive black holes in large haloes, though we defer a full investigation of Arkenstone in group and cluster-scale haloes to future work.
@@ 3.2.4 Galaxy sizes
Figure 7: Galaxy sizes at $z=0$, defined as the stellar half mass radii, as a function of stellar mass. Lines and points are displayed in the same way as Fig. 5, with observational data from Behroozi et al. (2022) included as a dotted line. Arkenstone tends to produce smaller low-mass galaxies, with the growth of larger galaxies happening at slightly smaller masses than in TNG.
As in previous studies that find that stellar winds have a significant impact on galaxy sizes, we find that the energy loading of our galactic winds plays a significant role in the stellar sizes of galaxies at $z=0$. In Fig. 7, we show the median 3D stellar half-mass radius as a function of galaxy stellar mass. TNG was calibrated to reproduce this relationship, though using a different observational dataset than the one presented here. The TNG runs (solid black line) slightly over-estimate the size of galaxies compared to the latest observational data (Behroozi et al., 2022), though we note that there are many caveats to such a comparison without full mock observations, and that MHD (as present in the original TNG simulations) tends to slightly reduce the size of stellar discs (Pillepich et al., 2018a). It is also worth noting that the TNG model significantly improved on the over-estimation of galaxy radii found in the original Illustris simulation, pointing again to the sensitivity of galaxy sizes to the feedback implementation (Genel et al., 2018).
With Arkenstone, the sizes of galaxies at a given stellar mass has a notable dependence on $\eta_{\mathrm{E}}$, with higher energy loadings leading to larger galaxies at nearly all galaxy masses (shown in the top panel of Fig. 7). The central value of $\eta_{\mathrm{E}}=0.3$ provides the best match to the observational galaxy sizes at the low mass end, though we note that the turning point where the slope steepens happens at lower masses in Arkenstone compared to TNG. This could be linked to the galaxy becoming quenched by the TNG BH model’s kinetic feedback, which kicks in at slightly lower masses in the Arkenstone runs due to larger BH masses. We postpone a more detailed study of the structure of galaxies to future work. The bottom two panels show that galaxy sizes are less affected by mass loading, and also change little when changing the slope of a variable energy loading. The variable loadings in general lie reasonably close to the observational estimates, although again a full forward modelling of galaxy sizes would be needed to draw more concrete conclusions.
@@ 3.3 CGM properties at $z=0$
Figure 9: Baryon fractions within $R_{200}$ at $z=0$, normalised by the universal baryon fraction ($f_{\mathrm{B}}=0.16$, Planck Collaboration et al., 2016). Lines and points are displayed in the same way as Fig. 5. Baryon fractions show a similar shape to gas fractions, increasingly gradually with halo mass until the kinetic mode of BH feedback begins to act.
@@
Figure 10: Mass-weighted mean circumgalactic medium temperatures within $R_{200}$ at $z=0$, normalised by the nominal virial temperature $T_{200}$. Lines and points are displayed in the same way as in Fig. 5. CGM temperatures as a function of halo mass tend to have a “U”-shape, with a minimum temperature (relative to $T_{200}$) at halo masses of $\sim\!10^{11-11.5}\,$M. Temperatures in Arkenstone runs tend to lie above those of TNG at $z=0$.
@@
Figure 11: Galaxy SFR as a function of stellar mass at $z=0$ for TNG (top) and the intermediate variable energy loading Arkenstone run ($\eta_{\rm E}\propto M_{\rm h}^{-1/3}$, bottom), coloured by the mass-weighted average gas fraction in the CGM (between twice the stellar half mass radius and $R_{200}$), normalised by the universal baryon fraction. The separated axes at the bottom of each panel show the galaxies with a zero SFR. Observational comparison data is from two sources: Salim et al. (2018), where the median is shown by a dashed black line and the shaded 1-$\sigma$ region has a backwards hash, and Alsing et al. (2024), which has a dotted median line and a forward hashed 1-$\sigma$ range. With Arkenstone, galaxies tend to have a slightly higher SFR than TNG, but a more suppressed gas fraction in the CGM – suggesting more preventative feedback.
@@
In Fig. 8 we show the median gas fraction within $R_{200}$, relative to the cosmic baryon fraction, as a function of halo mass. The shape of the TNG line is set by the combination of the stellar and BH feedback models. TNG was also calibrated to this quantity, though data exists only for groups (of which there are only a handful in our box) and clusters (of which there are none). As this paper focuses on haloes below this scale, which lies in a different part of parameter space to available observations, we do not show comparison data here.
@@
Below a halo mass of $\sim\!2\times 10^{12}$ M, the amount of gas in the CGM is primarily regulated by stellar feedback. We have already discussed how TNG tends to rely on ejection of mass from the ISM to regulate star formation. We can see how this ejection also removes much of the CGM at the lowest halo masses, while slightly larger haloes can retain more material. Above halo masses of $\sim\!2\times 10^{12}$ M, the kinetic mode of TNG’s AGN feedback typically kicks in, expelling much of the inner CGM. As the halo mass increases further, $R_{200}$ increases and haloes can gravitationally hold onto more of that material, leading to an increase in the gas fraction once more.
@@
A number of differences are present between runs with different energy loadings (top panel), with opposing trends at low and high halo masses. At low masses, the higher energy loadings lead to more preventative feedback, lowering the gas fraction that ends up within $R_{200}$. At high masses, the highest energy loading actually leads to the retention of more gas. This is likely due to the stunted growth of black holes (see Fig. 6) leading to less feedback, particularly because fewer BHs will then enter TNG’s powerful, BH mass-dependent kinetic mode.
The mass loading of Arkenstone winds makes little difference to the gas fraction of intermediate and high mass haloes for $\eta_{\mathrm{M}}\leq 0.3$. In low mass haloes the mass loading makes more of a difference, once again with lower mass loading leading to more preventative feedback, this time shown by the reduction of the CGM gas fraction for $\eta_{\mathrm{M}}=0.1$. The $\eta_{\mathrm{M}}=0.9$ run has particularly low gas fractions of 10 per cent or below in haloes with masses above $10^{12}$ M, at least a factor of two below the other runs, though as we show below, this is due to the gaseous halo being converted into (excess) stars.
The winds with variable energy loading smooth the abrupt change in median gas fraction present in TNG at halo masses of $\sim\!10^{10.5}$ M. Gas fractions within $R_{200}$ at $z=0$ in Arkenstone tend to lie below those of TNG – this CGM depletion can be seen in the regions close to galaxies in Fig. 3.
In Fig. 9 we show a similar plot, except now combining both the gas fractions and stellar fractions into the median baryon fraction. Here we see many of the same trends present in the SMHM for haloes below $10^{12}$ M, with the normalisation and slope of the relation strongly dependent on energy loading, and the mass loading affecting the normalisation. The baryon fraction of the $\eta_{\mathrm{M}}=0.9$ run at the massive end is very similar to all other runs, despite the significant difference in gas fraction - this run forms far too many stars. The median baryon fractions for the variable energy loading Arkenstone runs are very similar to TNG, with only slight differences between the different energy loading slopes.
@@
Preventative feedback can also act through the heating of the CGM (and IGM) around galaxies, inhibiting the cooling of gas onto the central galaxy. To study this, in Fig. 10 we show the mass-weighted temperature of all non-star forming gas cells (those not subject to the eEoS described in Section 2) within $R_{200}$, normalised by the halo virial temperature $T_{200}={\mu m_{\mathrm{p}}GM_{200}}/{(2k_{\mathrm{B}}R_{200})}$, where $\mu=0.59$ is the mean molecular weight of ionised primordial gas.
@@
The vast majority of haloes in all of the runs in Fig. 10 lie below 1, indicating a sub-virial mean temperature. The CGM temperatures in TNG are closest to the virial temperature at the lowest and highest masses, with a dip in between at $\sim\!10^{11-11.5}\,$M. This dip likely corresponds to haloes with a virial temperature near the peak of the cooling curve, leading to more efficient cooling (Fielding et al., 2017), particularly in TNG where the CGM density tend to be higher than in Arkenstone (see Fig. 8). Like the gas fraction plot, the relationship at low masses changes more smoothly in Arkenstone compared to TNG. The Arkenstone runs have a somewhat similar shape, with most runs having a minimum in CGM temperature in the halo mass range $\sim\!10^{11-11.5}\,$M, but with the average $z=0$ temperature being higher than TNG. There are no clear changes between the different Arkenstone simulations, with the exception of hotter CGM temperatures for the lowest mass galaxies when the wind’s specific energy is increased. We note that this median value of the mass-weighted mean temperatures of haloes hides significant complexity in the CGM, which generally cannot be characterised by a single temperature (Lochhaas et al., 2021).
@@
In Fig. 11 we see how this change in CGM properties drives the regulation of star formation. We show two-dimensional histograms of the SFR of galaxies as a function of stellar mass, highlighting the star-forming main sequence for TNG (top panel) and the variable energy loading Arkenstone run with $\eta_{\mathrm{E}}\propto M_{\mathrm{h}}^{-1/3}$ (bottom panel). For comparison we also show two observationally inferred main sequences, shown with dotted and dashed black lines (Salim et al., 2018; Alsing et al., 2024; Sommovigo et al., 2025). The histogram is coloured by the stellar mass-weighted average CGM gas fraction (defined as the gas within $R_{200}$ minus the gas within twice the stellar half mass radius), normalised by the cosmic baryon fraction. The darker regions therefore indicate where the CGM is relatively depleted of gas, which for Arkenstone typically occurs in galaxies below the star forming main sequence. TNG tends to have larger CGM gas fractions in general, as also shown in Fig. 8. It also tends to have slightly lower SFRs compared to Arkenstone  though the scatter in the observationally estimated star-forming main sequence precludes a stronger conclusion (see also Donnari et al., 2019).
3.4 Preventative feedback in action
Figure 12: Median mass outflow (top row) and inflow (bottom row) rates as a function of radius, binned in three different halo mass bins. The left and right columns show results at $z=0$ and $z=2$, respectively, with TNG shown in black and the Arkenstone variable energy loading run with $\eta_{\mathrm{E}}\propto M^{-1/3}$ in orange. The mass fluxes in Arkenstone are almost always lower than that of TNG, with the gap widening at smaller radii.
@@ 3.4 Preventative feedback in action
In this section we examine the degree to which feedback is ejective vs. preventative in Arkenstone simulations, by looking at mass fluxes directly, but also through investigating a subset of the quantities discussed in Sections 3.2 and 3.3 at $z=2$. Here, the star formation rate is higher and so the subsequent feedback is stronger. This allows us to illuminate the differences between how Arkenstone and TNG regulate star formation, even though they both can lead to similar $z=0$ galaxy properties.
Fig. 12 shows the total radial outflowing (top row) and inflowing (bottom row) gas mass flow rates as a function of radius, at $z=0$ (left column) and $z=2$ (right column). We calculate the mass fluxes at each radius $r$ by identifying the closest cell to each of $N_{\mathrm{pix}}=786,432$ healpix pixels (Górski et al., 2005) and assigning its properties to each pixel. The total mass flow rates then follow as
$\dot{M}=\frac{4\pi r^{2}}{N_{\mathrm{pix}}}\sum\rho v_{\mathrm{r}},$ (16)
which we split into outflow (where $v_{\mathrm{r}}>0$) and inflow rates (where $v_{\mathrm{r}}<0$). We do this for every central galaxy in each of three mass bins, and present the median value.
Turning first to the outflow rates, we can see how Arkenstone has consistently lower mass outflow rates than TNG at both $z=0$ and $z=2$, across all halo masses and radii. The difference between the two is notably smaller than the difference in input loadings (see Fig. 1) – the Arkenstone winds sweep up a significant mass of material as they move into and through the CGM.
The inflow rates are particularly interesting in the context of preventative feedback. First at $z=2$, where the global SFR is higher and feedback more active, we see that the median mass inflow rates in Arkenstone are consistently below those of TNG. The Arkenstone runs have the characteristic signature of preventative feedback – lower inflowing mass fluxes at smaller radii – showing how gas is prevented from moving towards the galaxy. This is particularly pronounced in the lower mass bin shown in Fig. 12, with a decrease of a factor 454-5 between $0.1R_{\mathrm{200}}$ and $2R_{\mathrm{200}}$. In contrast, the mass inflow in TNG remains largely flat as a function of radius, such that the inflowing mass flux is about 4 times higher at $0.1R_{\mathrm{200}}$ compared to Arkenstone. A similar trend is seen in the EAGLE simulations, which also has consistently lower inflow rates than TNG (Wright et al., 2024).
At $z=0$ this signature is less clear. We note that the magnitude of the inflowing (and outflowing) fluxes drops between $z=2$ and $z=0$, particularly due to both a lower average SFR and the ejective effect of AGN feedback in TNG’s BH model. Nevertheless, Arkenstone mass fluxes remain below TNG in all mass bins and at all radii, but often with a narrow gap than at $z=2$. In this Arkenstone run, we show in Fig. 11 that $z=0$ SFRs lie a little higher than TNG, which could potentially be linked to the narrower difference between Arkenstone and TNG in Fig. 12 at $z=0$.
@@
In Fig. 13 we show the SMHM at $z=2$, which shows mostly similar trends to $z=0$. Changing the energy loading of the hot wind has the greatest effect on the normalisation of the SMHM (top panel). The mass loading affects the normalisation and the slope of the SMHM at intermediate halo masses (central panel), with higher mass loadings leading to higher stellar masses. As discussed earlier, this relationship is due to lower mass loading causing a hotter and faster wind (with the same total power) and thus increased preventative feedback. The variable energy loadings (bottom panel) counter the overproduction of stars at the low mass end, bringing the SMHM closer to the observational estimates. Notably, most of the simulations lie above the observationally inferred SMHM at $z=2$ from Shuntov et al. (2022), though the Arkenstone variable energy loading runs are close.
@@
The most significant differences between the runs, which hold the key to understanding how Arkenstone differs from TNG, are shown in the CGM gas fraction (Fig. 14) and temperature (Fig. 15) plots.
@@
Turning first to the gas fractions, for TNG this starts off low at the smallest halo masses, before increasing to $\sim\!80$ per cent of the cosmic baryon fraction at a halo mass of $10^{12}$ M, and slowly declining beyond that as the black hole feedback takes effect. For Arkenstone  the gas fraction increase has a much shallower slope across halo masses. For simulations with different constant energy loading values (top panel), the normalisation for haloes below $10^{12}$ M is inversely proportional to $\eta_{\mathrm{E}}$. At the TNG ‘turnover’ mass of $10^{12}$ M, the Arkenstone runs have a gas fraction of 30-50 per cent of baryons, around half of TNG’s value at this point. The strong implication of this is a reduction in halo-scale accretion rates - either due to high specific energy winds pushing gas outside $R_{200}$, or the prevention of gas accreting within $R_{200}$ in the first place. This ‘prevention-via-CGM-depletion’ scenario, in contrast to ISM ejection, is a central feature of the Arkenstone model. Notably, the relationship between energy loading and gas fraction reverses in haloes above $10^{12}$ M, though this is likely due to the variations in stellar feedback affecting black hole growth and feedback.
@@
Varying the mass loading also affects the gas fractions (central panel) - the lowest mass loading of $\eta_{\mathrm{M}}=0.1$ also has the lowest gas fractions for low and intermediate mass galaxies, indicating that hotter and faster winds reduce the baryon content of the CGM even more at higher redshift. The variable energy loadings on the other hand lead to a gradually increasing gas fraction as a function of halo mass, up to a halo mass of $\sim\!10^{12}$ M. This slope is shallower than in TNG, and the median gas fraction values are below TNG at all halo masses.
@@
The temperature of the CGM is also an indicator of preventative feedback, with hotter, diffuse gas having longer cooling times and so reducing inflows onto the central galaxy. The CGM temperature at $z=2$ is shown in Fig. 15, defined in the same way as Fig. 10. Here we again see a strong difference between TNG and Arkenstone. The CGM temperatures for haloes around $\sim\!10^{11}\,$M are much cooler in TNG than in most Arkenstone runs, and the Arkenstone runs stay hotter throughout the entire halo mass range.
@@
The temperature varies with both energy and mass loading - higher specific energy winds lead to a hotter and more diffuse CGM, more able to prevent gas accreting onto galaxies and forming stars. The highest mass loaded Arkenstone run leads to a cooler CGM closer to TNG, as more energy is used to transport a heavier wind. The slope of the variable energy loading does not strongly impact the CGM temperature at $z=2$.
Across the parameter space studied in this work, the CGM of galaxies in Arkenstone is hotter and more diffuse, an indicator of the large role preventative feedback plays in regulating star formation.